OpenStudy (anonymous):

9x-4+3x How do I combine like terms?

OpenStudy (anonymous):

solve for x's so add the same variables, it's like normal addition.

OpenStudy (anonymous):

9x+3x = 4 first step

OpenStudy (anonymous):

12x=4

OpenStudy (anonymous):

x = 4/12

OpenStudy (anonymous):

x=1/3

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome